# The Monge optimal transport barycenter problem

### v1

@@ 9.6 Hidden patterns in ground-level atmospheric temperature
Switching to applications to real data, we consider next the hourly ground-level temperature in Ithaca, NY from 2007 to 2023. The data, available from National Oceanic and Atmospheric Administration, is displayed in the top panel of Figure 16. We will use the OTBP to investigate the dependence of this temperature on the diurnal and seasonal cycles, and to uncover hidden signals at the synoptic weather and multi-year scales.
We first solve the OTBP for $\rho(x|z_{1})$, where $x$ is the hourly temperature in Ithaca and $z_{1}\in[0,365.25]$ is the day of the year, a continuous, periodic factor. Panel (a) of Figure 14 displays the corresponding day-dependent median value of the simulated temperature $X(:,z_{1})$ and the corresponding conditional 90% confidence interval, capturing seasonal effects, superimposed for reference on the true observed temperatures for the year 2007. We use the conditional median and confidence intervals rather than the conditional mean and standard deviation of $x$ because they are more robust statistics and they are also much cheaper to compute: while computing the mean involves averaging $X(y,z)$ over all $\{y_{i}\}$ for each value of $z$, and similarly for the variance, the monotonicity of $X$ implies that the conditional median of $x$ is just $X(\bar{y},z)$, where $\bar{y}$ is the median of the $\{y_{i}\}$), and similarly for confidence intervals, since conditional percentiles of $\rho(x|z)$ translate directly from the corresponding percentiles in $y$. Next we consider instead $\rho(x|z_{2})$, where $z_{2}\in[0,24]$ is the time of the day, another continuous and periodic factor, displaying in panel (b) the simulated median diurnal cycle and corresponding confidence interval together with the true $x$ for 2007. Then we combine the two factors and consider $\rho(x|z_{1},z_{2})$, with results displayed on panel (c) both for the the full year 2007 at once and for zoomed-in versions for each season. One may notice in all panels how the 90% confidence interval depends on $z$, often adopting asymmetric shapes around the median and displaying for instance interesting contrasts between day and night. This is one manifestation of the power of having captured the full conditional distribution $\rho(x|z)$, as opposed to just a few statistics, such as the conditional mean value computed in regression.
@@
(a) Using only day of year as a factor
(b) Using only hour of day as a factor
(c) Using the two periodic time factors together
@@
We switch next to consider the variability of $x$ not explained by $(z_{1},z_{2})$, as captured in $y$. In order to analyze both synoptic weather and short-term multi-year variability, we introduce two new time factors, $z_{3}$ and $z_{4}$, built by rescaling time using different scales. In other words, both $z_{3}$ and $z_{4}$ consist just of the time $t$ (measured in hours), but the bandwidths used for the corresponding kernels are of the order of 30 and 3 days, respectively (we set these scales through the parameter $\gamma_{z}$ defined in the appendix.) As we did for $z_{1,2}$, we remove the variability in $y$ attributable to $z_{3}$ and $z_{4}$ separately, then together. Figures 16 and 17 show the results of introducing these new factors. Panels (a) and (b) of figure 16 display the original time series $x(t)$, the signal $y_{1}(t)$ resulting from removing $z_{1,2}$ from $x$ and the signal $y_{2}(t)$ resulting from removing $z_{3}$ and $z_{4}$ from $y_{1}$. The removal of variability is reflected in the signals’ decreasing variance, from $113.71$ for $x$, through $24.71$ for $y_{1}$, to $17.10$ for $y_{2}$. Beyond the decreasing variance, one can observe the further explanation of variability in the fact that $y_{2}$ is much more homogeneous in time than $y_{1}$, which has a clear inhomogeneity associated with the synoptic weather signal (The fact that similarly $y_{1}$ does not display the time dependence on the seasonal cycle present on $x$ is far more obvious to the eye.) Panel (a) of figure 17 displays the median temperature dependence on $z_{3}$, a multi-year signal, and panel (b) the dependence on $z_{4}$, corresponding to synoptic weather, over the year 2007. The median temperature and 90% confidence interval determined by $z_{3}$ and $z_{4}$ together is displayed in panel (c) and zoomed-in over 2007 in panel (d). A climate scientist looking at these reconstructions may not only confirm that the method has captured the right scales (a roughly 2-4 year scale for the multiyear signal and around 15 days for the synoptic weather) but also detect individual signals, such as in panel (a) a signal resulting from the El Niño years 2007, 2010 and 2016, and in panel (b) a signal from the North American heat wave of 2007, which may have contributed to the elevated temperature of Ithaca during the late summer and early autumn. Finally, panel (e) displays the full $z_{1,2,3,4}$-dependent conditional median $\bar{x}(z(t))$ and 90% confidence interval over 2007. We can see in this plot not only how well the reconstruction has captured the dependence of temperature on time, season and the synoptic and multi-year time scales, but also how it has not captured (by construction) weather signals shorter than a week long. These could of course be captured by another factor $z_{5}$ with shorter bandwidths, but doing this would take us too far afield in the context of the current article. Notice that these shorter scales are nonetheless represented as noise in the 90 percentile, a general property of the OTBP methodology: as new factors $z^{l}$ are introduced, these explain away part of the variability previously present in the conditional distribution $\rho(x|z)$.
@@
(a) The original time series $x(t)$, the $y_{1}(t)$ resulting from removing the effects of periodic time factors (time of day and day of year), and the $y_{2}(t)$ resulting from further removing from $y_{1}$ the synoptic weather and multi-year signals
(b) A zoom-in version of the previous figure, restricted to 2007
Figure 16: Effect of removing from the temperature $x$, in succession, the effect of periodic time factors (time of day and day of year) and of time itself at the synoptic and multiyear scales.
(a) Median multi-year trend
(b) Median synoptic weather trend (2007)
(c) Median temperature and 90% confidence interval predicted by $z_{3}$ and $z_{4}$ together
(d) Zoom of the above for 2007 superposed on $y_{1}(t)$
(e) Median temperature and 90% confidence interval as a functions of day of year, time of day, multi-year trend and the synoptic weather trend, displayed over the true temperature for 2007
@@ 9.7 Forecasting of global ocean states
We further illustrate the methodology, using it to forecast six months ahead the global sea surface temperature (SST). The data (available at Met Office Hadley Centre observations) consists of monthly values of the SST from 1870 to 2024, over a global $1\times 1$ latitude-longitude grid. The resulting dataset $T_{i,j}^{l}$ has a dimension of $180\times 360\times 1860$, corresponding to latitude (indexed by $i$), longitude ($j$) and time ($l$ in months). The goal is to use historical observations to predict the global SST 6 months into the future.
In order to extract a lower dimensional time signal from the data, we apply a standard pre-processing to the whole dataset:
  • Filter out spatial grid points that either lie over land (where SST is undefined) or contain missing data, resulting in $31{,}094$ valid spatial grid points; all subsequent analysis, including EOF computation, is restricted to this filtered spatial domain;
@@
(We could replace the standard pre-processing by a far more informative one based on the OTBP methodology, but this would take us too far afield in the current article, it is currently investigated as part of a general methodology for the analysis of high-dimensional time series.) It is known [41] that the first EOF component correlates strongly with ENSO events. Figure 18 depicts the first 3 $\text{EOF}^{k}(x)$ as well as their temporal coefficients $\text{C}^{1,2,3}(t)$. The prediction task then reduces to forecasting the coefficients $\text{C}^{k}$ from their lagged observations $\text{C}^{l}(t-\Delta t)$, with $\Delta t=6$ months.
@@
(a) The first 3 coefficients $C^{k}(t)$ (black) and our prediction (red for prediction mean, and pink for one standard deviation away from prediction mean).
(b) Anomaly at 3 dates (from left to right: strong El Niño, neutral, strong La Niña), color limit: ±$\pm 1.5^{\circ}C$. The top row displays the ground truth, while the bottom row contains our prediction, truncated to the top 50 components.
@@ 10 Summary and discussion
We have developed in this article an efficient methodology for solving the sample-based Monge optimal transport barycenter problem, which takes as input $n$ observed sample pairs $\{x_{i},z_{i}\}$ drawn from an unknown joint distribution $\pi(x,z)$, and produces as output $n$ associated samples $\{y_{i}\}$ from the barycenter $\mu$ of the conditional distributions $\rho(x|z)$ under $\gamma(z)=\pi(\mathcal{X},z)$. In addition, it produces $n$ samples $\{x_{i}^{*}\}$ drawn from the estimated $\rho(:|z_{*})$ for any proposed target value $z_{*}$ of the covariates $z$ and it estimates $\rho(x|z)$, instrumental for model-free Bayesian inference. A corollary extends the procedure to solve the regular [Monge] optimal transport problem.
Central to the methodology and to most of its applications is a formulation of the OTBP not in terms of the barycenter $\mu$ itself but of the underlying random variable $y=T(x,z)$, which must be statistically independent of the factors $z$. A test-based formulation of independence through the uncorrelation between all functions $\{f(z),g(y)\}$ within suitable functional spaces $\{\mathcal{F},\mathcal{G}\}$ provides an adversarial formulation of the OTBP. Since the best adversarial functions $f$ and $g$ can be found exactly in terms of the first principal components of a matrix $A(y)$, the problem reduces to a single minimization over the map $T$. Solving this problem through gradient descent over $y$ yields a flow in phase space that transports each $y_{i}$ from $x_{i}$ to $T(x_{i},z_{i})$. Fortiutously, the resulting map $T$ can be inverted in closed form, which facilitates much both the simulation and the estimation of $\rho(x|z)$. A byproduct of this closed-form inversion is the extraction of factors $\{f^{k}\}(z)$ that encode the dependence of $x$ on $z$.
@@
11 Appendix: a data-adapted functional space
@@
11.1 Embedding $z$ in an Euclidean space
Since the components of $z$ can be of arbitrary type, including real, periodic, categorical and more (no so those of $y$, which are typically real), we first embed $z$ in an Euclidean space $R^{k}$ as follows. For each component $z^{l}\not\in R$ of $z$,
  1. 1.
    When $z^{l}$ is periodic with period $T$, we embed it in $R^{2}$, mapping $z^{l}$ to $w$ on the unit circle,
    $w\left(z^{l}\right)=\left[\cos\left(\frac{2\pi}{T}z^{l}\right),\sin\left(\frac{2\pi}{T}z^{l}\right)\right].$
  2. 2.
    When $z^{l}$ is categorical with $h$ discrete values $v_{k}$, we embed it in $R^{h-1}$, mapping the $\{v_{k}\}$ to $h$ equidistant points, the vertices of a regular simplex.
  3. 3.
    For variables $z$ of a more complex type, such as images, distributions or graphs, we introduce an application-specific distance among them and embed them into some $R^{h}$ accordingly.
Having done this, we can restrict attention to $\mathcal{Z}=R^{d_{z}}$.
11.2 Approximation of functions through mollifiers of the delta function
@@
[…]
(Notice that the last expression is only valid for values of $z$ within the support of the distribution $\gamma$.)
Mollifying $\delta(x-y)$ into a smooth non-negative function $K(x,y)$ that concentrates near $x=y$ and integrates to $1$ over $x$, we have
[…]
which in terms of samples $z^{c}_{j}\sim\gamma(z)$ yields the empirical version
@@
Notice that the normalization factor for the weights in this formula is different from the one in kernel regression, which is $z$-dependent:
$w^{j}_{KR}=\frac{K\left(z,z^{c}_{j}\right)}{\sum_{h}K\left(z,z^{c}_{h}\right)}.$
Unlike kernel regression, our approximation is based on the distribution underlying the $\{z^{c}_{j}\}$. Notice also that nothing in our argument requires $K(x,y)$ to be a symmetric function, so it is not a “kernel” in the conventional sense. This allows us to use center-dependent bandwidths, coarser where the data is sparser.
If $K$ is a smooth function of $z$, so is the right-hand side of (19) for any choice of $f\left(z^{c}_{j}\right)$. We conclude that the expression
$f(z)=\sum_{j}a_{j}\frac{K\left(z,z^{c}_{j}\right)}{\sum_{h}K\left(z^{c}_{j},z^{c}_{h}\right)}$
parameterizes arbitrary smooth functions of $z$ within the support of $\gamma(z)$. Since the denominator does not depend on $z$, we can absorb it into the definition of $a_{j}$, which yields
@@
where it is no longer required that the $\{F^{j}(z)\}$ integrate to one, since the corresponding normalizing constants can also be absorbed into the $\{a_{j}\}$. Then the space $\mathcal{F}$ of smooth functions $f(z)$ agrees with the column space of the operator $F$. In order to consider functions with zero mean, it is enough to subtract the mean of each column of $F$.
@@
There is no need for the set of centers $\{z^{c}_{j}\}$ for $K$ and the set of points $\{z_{i}\}$ where $f$ is to be evaluated to agree; it is enough that the support of the distribution $\gamma$ underlying the former contains the support of the latter. Thus, when $z$ is restricted to the sample points $\{z_{i}\}$, $F(z)$ is a rectangular matrix. Using a number $m\ll n$ of centers reduces the computational cost associated to evaluating the kernels. Moreover, when applied to $g(y)$, one needs to decouple the centers $\{y^{c}_{j}\}$ from the samples $\{y_{i}\}$, as only the latter are arguments over which the objective function $L$ is minimized.
The centers $\{z^{c}_{i}\}$ should be well-balanced and representative of the distribution underlying the $\{z_{i}\}$. A simple procedure for selecting $m\ll n$ centers satisfying these conditions is through k-means applied to the $\{z_{i}\}$, which has been the choice adopted for all examples in Section 9, where we have set $m=\min([\sqrt{n}],m_{max})$, with $m_{max}$ set by the user. Notice that, when we embed periodic or discrete $z^{l}$ in $R^{d}$, the corresponding $l$-component of the centers $z^{c}_{j}$ need not lie on the unit circle or on the vertices of a simplex.
We adopted for $K(z,z^{c})$ a Gaussian function with center-dependent inverse covariance matrix $S_{j}=S({z^{c}_{j}})$:
$K\left(z,z^{c}_{j}\right)=e^{-\frac{1}{2}\left(z-z_{j}^{c}\right)^{\prime}S_{j}\left(z-z_{j}^{c}\right)}\ .$
11.3 Data-adapted determination of bandwidths
Tuning the $\{S_{j}\}$ is critical for extracting as much dependence of $x$ on $z$ as possible, by capturing the right functions $f(z)$ and $g(y)$. A simple example illustrates how the appropriateness of a functional space for $f(z)$ depends not just on the samples $\{z_{i}\}$ but also on the $\{x_{i}\}$. Consider a situation where $x$ depends only on the first component $z^{1}$ of a multidimensional $z\in R^{k}$, i.e. $x\sim\rho(:|z^{1})$. Capturing this dependence requires a family of functions $f(z)$ that depend only on $z^{1}$. The ideal $S_{j}$ for this adopt the form
$S_{j}\propto e_{1}{e_{1}}^{\prime},\quad e_{1}=(1,0,\ldots,0)^{\prime},$
since these yield functions $K(z,z^{c})$ of $z^{1}$ alone, disregarding all other components of both $z$ and $z^{c}$. One would not have been able to make these selection from only looking at the $\{z_{i}\}$.
Similarly, consider a situation where the distribution of $x$ depends on $z$ through some functions $f_{k}(z)$, with the strength of this dependence quantified by numbers $\sigma_{k}$, $0\leq\sigma_{k}\leq 1$. An appropriate set $\{S_{j}\}$ could adopt the form
$S_{j}\propto\sum_{k}\sigma_{k}v_{k}{v_{k}}^{\prime},\quad v_{k}=\nabla_{z}f_{k}(z)\big{|}_{z^{c}_{j}},$
so that $K(z,z^{c})$ changes only in the span of the local $\{\nabla_{z}f_{k}\}$, weighted by their relevance. Thus, an appropriate choice of the $\{S_{j}\}$ yields mollifiers $K_{j}(z,z^{c}_{j})$ of $\delta(z-z^{c}_{j})$ that single out the sub-manifold of the tangent space to $Z$ at $z^{c}_{j}$ on which $y$ may depend.
Yet we do not know the form of this dependence before hand, since determining it is precisely our algorithm’s goal. Thus we first make a choice based not on the relation between $x$ and $z$ but on the data available for each. We can subsequently refine this choice by iteratively capturing the dependence between $z$ and $x$, though we do not pursue such refinement within this article, other than through some straightforward cross-validation sketched below.
11.3.1 Initial determination of the $\{S^{0}_{j}\}$
The most natural function $f(z)$ to attempt to capture when looking only at the $\{z_{i}\}$ is their underlying probability density $\gamma(z)$. We apply the following adaptive procedure to determine the corresponding $\{S_{j}^{0}\}$.
  1. 1.
@@ 11.3.1 Initial determination of the $\{S^{0}_{j}\}$
  2. 2.
    Introducing an adjustable parameter $\alpha$, define
    $K_{i}^{j}=K_{g}\left(z^{c}_{i},z^{c}_{j}\right)=e^{-\frac{1}{2\alpha^{2}}\left(z^{c}_{i}-z^{c}_{j}\right)^{\prime}S_{g}\left(z^{c}_{i}-z^{c}_{j}\right)},$
    estimate $\gamma^{i}_{\alpha}\left(z^{c}_{i}\right)$ through leave-one-out kernel density estimation,
    $\gamma^{i}_{\alpha}\left(z^{c}_{i}\right)\propto\frac{1}{\alpha^{d}}\sum_{j\neq i}K_{i}^{j},$
    determine $\alpha$ through leave-one-out maximal likelihood,
    $\alpha^{*}=\arg\max_{\alpha}L=\sum_{i=1}^{m}\log\left(\gamma^{i}_{\alpha}\left(z^{c}_{i}\right)\right),$
    and define accordingly $S_{\alpha_{*}}=\frac{S_{g}}{\alpha_{*}^{2}}$. A practical choice is to maximize $L$ over a finite set of candidate $\alpha$’s centered around the rule-of-thumb value
    $\alpha_{r.o.th.}=\frac{\left(\frac{4}{d+2}\right)^{\frac{1}{d+4}}}{m^{\frac{1}{d+4}}}.$
  3. 3.
    Next rescale the $S_{\alpha_{*}}$ locally using the estimated $\gamma^{i}_{\alpha_{*}}$,
    $S_{j}=\left(\gamma_{\alpha_{*}}\left(z^{c}_{j}\right)\right)^{\frac{2}{d}}S_{\alpha_{*}},$
    so that the the number of points $\{z^{c}_{i}\}$ within the effective support of the corresponding function $K(z,z^{c}_{j})$ is roughly independent of $j$.
  4. 4.
    Finally, introducing a new global adjustable parameter $\beta$, write
    $K_{j}^{\beta}(z)=e^{-\frac{1}{2\beta^{2}}\left(z-z^{c}_{j}\right)^{\prime}S_{j}\left(z-z^{c}_{j}\right)},$
    with $\beta$ determined again through leave-one-out maximal likelihood:

### v2

@@ 9.6 Hidden patterns in ground-level atmospheric temperature
We consider next the hourly ground-level temperature in Ithaca, NY from 2007 to 2023. The data, available from National Oceanic and Atmospheric Administration, is displayed in the top panel of Figure 12. We will use the OTBP to investigate the dependence of this temperature on the diurnal and seasonal cycles, and to uncover hidden signals at the synoptic weather and multi-year scales.
We first solve the OTBP for $\rho(x|z_{1})$, where $x$ is the hourly temperature in Ithaca and $z_{1}\in[0,365.25]$ is the day of the year, a continuous, periodic factor. Panel (a) of Figure 10 displays the corresponding day-dependent median value of the simulated temperature $X(:,z_{1})$ and the corresponding conditional 90% confidence interval, capturing seasonal effects, superimposed for reference on the true observed temperatures for the year 2007. We use the conditional median and confidence intervals rather than the conditional mean and standard deviation of $x$ because they are more robust statistics and they are also much cheaper to compute: while computing the mean involves averaging $X(y,z)$ over all $\{y_{i}\}$ for each value of $z$, and similarly for the variance, the monotonicity of $X$ implies that the conditional median of $x$ is just $X(\bar{y},z)$, where $\bar{y}$ is the median of the $\{y_{i}\}$), and similarly for confidence intervals, since conditional percentiles of $\rho(x|z)$ translate directly from the corresponding percentiles in $y$. Next we consider instead $\rho(x|z_{2})$, where $z_{2}\in[0,24]$ is the time of the day, another continuous and periodic factor, displaying in panel (b) the simulated median diurnal cycle and corresponding confidence interval together with the true $x$ for 2007. Then we combine the two factors and consider $\rho(x|z_{1},z_{2})$, with results displayed on panel (c) for the the full year 2007. One may notice in all panels how the 90% confidence interval depends on $z$, often adopting asymmetric shapes around the median and displaying interesting contrasts between day and night. This is one manifestation of the power of capturing the full conditional distribution $\rho(x|z)$, as opposed to just a few statistics, such as the conditional mean value computed in regression.
@@
We switch next to consider the variability of $x$ not explained by $(z_{1},z_{2})$, as captured in $y$. In order to analyze both synoptic weather and short-term multi-year variability, we introduce two new time factors, $z_{3}$ and $z_{4}$, built by rescaling time using different scales. In other words, both $z_{3}$ and $z_{4}$ consist just of the time $t$ (measured in hours), but the bandwidths used for the corresponding kernels are of the order of 30 and 3 days, respectively (we set these scales through the parameter $\gamma_{z}$ defined in the appendix.) As we did for $z_{1,2}$, we remove the variability in $y$ attributable to $z_{3}$ and $z_{4}$ separately, then together. Figures 12 and 13 show the results of introducing these new factors. Panels (a) and (b) of figure 12 display the original time series $x(t)$, the signal $y_{1}(t)$ resulting from removing $z_{1,2}$ from $x$ and the signal $y_{2}(t)$ resulting from removing $z_{3}$ and $z_{4}$ from $y_{1}$. The removal of variability is reflected in the signals’ decreasing variance, from $113.71$ for $x$, through $24.71$ for $y_{1}$, to $17.10$ for $y_{2}$. Beyond the decreasing variance, one can observe the further explanation of variability in the fact that $y_{2}$ is much more homogeneous in time than $y_{1}$, which has a clear inhomogeneity associated with the synoptic weather signal (The fact that similarly $y_{1}$ does not display the time dependence on the seasonal cycle present on $x$ is far more obvious to the eye.) Panel (a) of figure 13 displays the median temperature dependence on $z_{3}$, a multi-year signal, and panel (b) the dependence on $z_{4}$, corresponding to synoptic weather, over the year 2007. The median temperature and 90% confidence interval determined by $z_{3}$ and $z_{4}$ together is displayed in panel (c) and zoomed-in over 2007 in panel (d). A climate scientist looking at these reconstructions may not only confirm that the method has captured the right scales (a roughly 2-4 year scale for the multiyear signal and around 15 days for the synoptic weather) but also detect individual signals, such as in panel (a) a signal resulting from the El Niño years 2007, 2010 and 2016, and in panel (b) a signal from the North American heat wave of 2007, which may have contributed to the elevated temperature of Ithaca during the late summer and early autumn. Finally, panel (e) displays the full $z_{1,2,3,4}$-dependent conditional median $\bar{x}(z(t))$ and 90% confidence interval over 2007. We can see in this plot not only how well the reconstruction has captured the dependence of temperature on time, season and the synoptic and multi-year time scales, but also how it has not captured (by construction) weather signals shorter than a week long. Notice that these shorter scales are nonetheless represented as noise in the 90 percentile, a general property of the OTBP methodology: as new factors $z^{l}$ are introduced, these explain away part of the variability previously present in the conditional distribution $\rho(x|z)$.
@@
Figure 12: The original time series $x(t)$, the $y_{1}(t)$ resulting from removing the effects of periodic time factors (time of day and day of year), and the $y_{2}(t)$ resulting from further removing from $y_{1}$ the synoptic weather and multi-year signals. On the right panels, a zoom-in version restricted to 2007.
@@ 9.7 Forecasting of global ocean states
We further illustrate the methodology, using it to forecast six months ahead the global sea surface temperature (SST). The data (available at Met Office Hadley Centre observations) consists of monthly values of the SST from 1870 to 2024, over a global $1\times 1$ latitude-longitude grid. The resulting dataset $T_{i,j}^{l}$ has a dimension of $180\times 360\times 1860$, corresponding to latitude (indexed by $i$), longitude ($j$) and time ($l$ in months). The goal is to use historical observations to predict the global SST 6 months into the future. In order to extract a lower dimensional time signal from the data, we apply a standard pre-processing to the whole dataset:
  • Filter out spatial grid points that either lie over land (where SST is undefined) or contain missing data, resulting in 31,09431{,}094 valid spatial grid points; all subsequent analysis, including EOF computation, is restricted to this filtered spatial domain;
@@
(We could replace the standard pre-processing by a far more informative one based on the OTBP methodology itself.) It is known [42] that the first EOF component correlates strongly with ENSO events. Figure 14 depicts the first 3 $\text{EOF}^{k}(x)$ as well as their temporal coefficients $\text{C}^{1,2,3}(t)$. The prediction task then reduces to forecasting the coefficients $\text{C}^{k}$ from their lagged observations $\text{C}^{l}(t-\Delta t)$, with $\Delta t=6$ months.
@@ 10 Summary and discussion
This article develops an efficient methodology for solving the sample-based Monge optimal transport barycenter problem, which takes as input $n$ observed sample pairs $\{x_{i},z_{i}\}$ drawn from an unknown joint distribution $\pi(x,z)$, and produces as output $n$ associated samples $\{y_{i}\}$ from the barycenter $\mu$ of the conditional distributions $\rho(x|z)$ under $\gamma(z)=\pi(\mathcal{X},z)$. In addition, it produces $n$ samples $\{x_{i}^{*}\}$ drawn from the estimated $\rho(:|z_{*})$ for any proposed target value $z_{*}$ of the covariates $z$ and it estimates $\rho(x|z)$, instrumental for model-free Bayesian inference. A corollary extends the procedure to solve the regular [Monge] optimal transport problem.
Central to the methodology and its applications is a formulation of the OTBP not in terms of the barycenter $\mu$ itself but of the underlying random variable $y=T(x,z)$, which must be statistically independent of the factors $z$. A test-based formulation of independence through the uncorrelation between all functions $\{f(z),g(y)\}$ within suitable functional spaces $\{\mathcal{F},\mathcal{G}\}$ provides an adversarial formulation of the OTBP. Since the best adversarial functions $f$ and $g$ can be found exactly in terms of the first principal components of a matrix $A(y)$, the problem reduces to a single minimization over the map $T$. Solving this problem through gradient descent over $y$ yields a flow that transports each $y_{i}$ from $x_{i}$ to $T(x_{i},z_{i})$. The resulting map $T$ can be inverted in closed form, which facilitates both the simulation and the estimation of $\rho(x|z)$. A byproduct of this closed-form inversion is the extraction of factors $\{f^{k}\}(z)$ that encode the dependence of $x$ on $z$.
@@
Appendix A Appendix: a data-adapted functional space
@@
Since the components of $z$ can be of arbitrary type, including real, periodic, categorical and more (not so those of $y$, which are typically real), we first embed $z$ in an Euclidean space $R^{k}$ as follows. For each component $z^{l}\not\in R$ of $z$: 1) when $z^{l}$ is periodic with period $T$, we embed it in $R^{2}$, mapping $z^{l}$ to $w$ on the unit circle, $w\left(z^{l}\right)=\left[\cos\left(\frac{2\pi}{T}z^{l}\right),\sin\left(\frac{2\pi}{T}z^{l}\right)\right]$; 2) when $z^{l}$ is categorical with $h$ discrete values $v_{k}$, we embed it in $R^{h-1}$, mapping the $\{v_{k}\}$ to $h$ equidistant points, the vertices of a regular simplex; 3) for variables $z$ of a more complex type, such as images, distributions or graphs, we introduce an application-specific distance among them and embed them into some $R^{h}$ accordingly. Having done this, we can restrict attention to $\mathcal{Z}=R^{d_{z}}$.
@@
$\gamma(z)=\int\gamma\left(z^{\prime}\right)\delta\left(z-z^{\prime}\right)dz^{\prime},\quad f(z)=\int f\left(z^{\prime}\right)\delta\left(z-z^{\prime}\right)dz^{\prime}=\int f\left(z^{\prime}\right)\frac{\delta\left(z-z^{\prime}\right)}{\int\gamma\left(z^{\prime\prime}\right)\delta\left(z^{\prime}-z^{\prime\prime}\right)dz^{\prime\prime}}\ \gamma\left(z^{\prime}\right)dz^{\prime}$
(Notice that the last expression is only valid for values of $z$ within the support of the distribution $\gamma$.) Mollifying $\delta(x-y)$ to a non-negative function $K(x,y)$ that concentrates near $x=y$ and integrates to $1$ over $x$ yields $f(z)\approx\int f\left(z^{\prime}\right)\frac{K\left(z,z^{\prime}\right)}{\int\gamma\left(z^{\prime\prime}\right)K\left(z^{\prime},z^{\prime\prime}\right)dz^{\prime\prime}}\gamma\left(z^{\prime}\right)dz^{\prime}$, which in terms of samples $z^{c}_{j}\sim\gamma(z)$ yields the empirical version
@@
Notice that nothing in our argument requires $K(x,y)$ to be a symmetric function, so it is not a “kernel” in the conventional sense. This allows us to use center-dependent bandwidths, coarser where the data is sparser.
If $K$ is a smooth function of $z$, so is the right-hand side of (18) for any choice of $f\left(z^{c}_{j}\right)$. We conclude that $f(z)=\sum_{j}a_{j}\frac{K\left(z,z^{c}_{j}\right)}{\sum_{h}K\left(z^{c}_{j},z^{c}_{h}\right)}$ parameterizes arbitrary smooth functions of $z$ within the support of $\gamma(z)$. Since the denominator does not depend on $z$, we can absorb it into the definition of $a_{j}$, which yields
@@
where it is no longer required that the $\{F^{j}(z)\}$ integrate to one, since the corresponding normalizing constants can also be absorbed into the $\{a_{j}\}$. Then the space $\mathcal{F}$ of smooth functions $f(z)$ agrees with the column space of the operator $F$. To consider functions with zero mean, it is enough to subtract the mean of each column of $F$.
@@
There is no need for the set of centers $\{z^{c}_{j}\}$ for $K$ and the set of points $\{z_{i}\}$ where $f$ is to be evaluated to agree; it is enough that the support of the distribution $\gamma$ underlying the former contains the support of the latter. Using a number $m\ll n$ of centers reduces the computational cost associated to evaluating the kernels. Moreover, when applied to $g(y)$, one needs to decouple the centers $\{y^{c}_{j}\}$ from the samples $\{y_{i}\}$, as only the latter are arguments over which the objective function $L$ is minimized. A simple procedure for selecting $m\ll n$ centers is through k-means applied to the $\{z_{i}\}$, which has been the choice adopted for all examples in Section 9, where we have set $m=\min([\sqrt{n}],m_{max})$, with $m_{max}$ set by the user. We adopted for $K(z,z^{c})$ a Gaussian function with center-dependent inverse covariance matrix $S_{j}=S({z^{c}_{j}})$: $K\left(z,z^{c}_{j}\right)=e^{-\frac{1}{2}\left(z-z_{j}^{c}\right)^{\prime}S_{j}\left(z-z_{j}^{c}\right)}$.
Tuning the $\{S_{j}\}$ is critical for extracting as much dependence of $x$ on $z$ as possible, by capturing the right functions $f(z)$ and $g(y)$. The appropriateness of a functional space for $f(z)$ depends not just on the samples $\{z_{i}\}$ but also on the $\{x_{i}\}$. Yet we do not know the form of this dependence before hand, since determining it is precisely our algorithm’s goal. Thus we make a choice based not on the relation between $x$ and $z$ but on the data available for each. The most natural function $f(z)$ to attempt to capture when looking only at the $\{z_{i}\}$ is their underlying probability density $\gamma(z)$. We apply the following adaptive procedure to determine the corresponding $\{S_{j}^{0}\}$.
@@
Introducing an adjustable parameter $\alpha$, define $K_{i}^{j}=K_{g}\left(z^{c}_{i},z^{c}_{j}\right)=e^{-\frac{1}{2\alpha^{2}}\left(z^{c}_{i}-z^{c}_{j}\right)^{\prime}S_{g}\left(z^{c}_{i}-z^{c}_{j}\right)}$, estimate $\gamma^{i}_{\alpha}\left(z^{c}_{i}\right)$ through leave-one-out kernel density estimation, $\gamma^{i}_{\alpha}\left(z^{c}_{i}\right)\propto\frac{1}{\alpha^{d}}\sum_{j\neq i}K_{i}^{j}$, determine $\alpha$ through leave-one-out maximal likelihood, $\alpha^{*}=\arg\max_{\alpha}L=\sum_{i=1}^{m}\log\left(\gamma^{i}_{\alpha}\left(z^{c}_{i}\right)\right)$, and define accordingly $S_{\alpha_{*}}=\frac{S_{g}}{\alpha_{*}^{2}}$. A practical choice is to maximize $L$ over a finite set of candidate $\alpha$’s centered around the rule-of-thumb value $\alpha_{r.o.th.}=\frac{\left(\frac{4}{d+2}\right)^{\frac{1}{d+4}}}{m^{\frac{1}{d+4}}}$. Rescale the $S_{\alpha_{*}}$ locally using the estimated $\gamma^{i}_{\alpha_{*}}$, $S_{j}=\left(\gamma_{\alpha_{*}}\left(z^{c}_{j}\right)\right)^{\frac{2}{d}}S_{\alpha_{*}}$, so that the the number of points $\{z^{c}_{i}\}$ within the effective support of the corresponding function $K(z,z^{c}_{j})$ is roughly independent of $j$. Finally, introducing a new global adjustable parameter $\beta$, write $K_{j}^{\beta}(z)=e^{-\frac{1}{2\beta^{2}}\left(z-z^{c}_{j}\right)^{\prime}S_{j}\left(z-z^{c}_{j}\right)}$, with $\beta$ determined again through leave-one-out maximal likelihood, and define $S^{0}_{j}=\frac{S_{j}}{\beta_{*}^{2}}$: